Two functions f and g are defined on the set R of real numbers by
f(x) = ax^2 + bx − 2 and g(x) = 3x + 2. If f(1) = 7 and f(2) = 10. Find:
(a) the values of a and b;
(b) the value(s) of x if f(x) = g(x) − 1.
1
Expert's answer
2020-07-12T17:31:34-0400

(a) "f(1)=7" and "f(2)=10" give "a+b-2=7" and "4a+2b-2=10". By the first equation, "b=9-a", so substituting this to the second one gives

"4a+2(9\u2212a)=12\u21d2a+9=6\u21d2a=\u22123,b=12".

(b) We are to solve the quadratic equation

"-3x^2+12x-2=3x+1\Leftrightarrow 0=3x^2-9x+3\Rightarrow x=\frac{3\pm\sqrt5}{2}"